# Learning to Predict Program Execution by Modeling Dynamic Dependency on Code Graphs

### v2

@@ XI-A Data Processing and Evaluation Metrics
In this section, we focus on a new metric, Error Localization Accuracy, which measures the number of times the predicted buggy line matches the actual buggy line. Based on the results from Section X, we observed that Claude performed the best in detecting runtime errors. Therefore, for this experiment, we use Claude as the main baseline to compare with our model.
@@ XI-B1 HeatMap Visualization
Alpha Value Error Localization Acc
$\alpha=0.5$ 72.22%
$\alpha=0.7$ 74.82%
$\alpha=0.9$ 77.40%
$\alpha=0.95$ 78.31%
@@ XII-A Fuzz Testing Procedure
  • Input Generation: We employed the LLM Claude to generate inputs that are likely to raise runtime errors in the provided code snippets.
  • Runtime Error Detection: The code snippet with the generated inputs encoded in the assignment statements at the beginning was fed into CodeFlow. The model processed the code to determine whether it contained a runtime error and, if so, localized the specific buggy statement.
@@ XII-B Empirical Results
Error Localization Acc 42.27% 49.53% 50.00% 42.61%
To evaluate the effectiveness of CodeFlow in supporting fuzz testing, we conducted experiments on 50 buggy Python code snippets from the FixEval dataset in which we removed all input variables and import statements. The results, summarized in Table VI, show that while Claude by itself detected 32 runtime errors, incorporating its fuzzing with CodeFlow detected 44 errors within 30 seconds (37.5% relative improvement), 46 errors within 60 seconds, and 47 errors within 120 seconds.
Additionally, the Error Localization Accuracy with CodeFlow was consistently higher than that of Claude, as demonstrated in RQ3. It is particularly challenging for LLMs like Claude to detect the correct line containing bugs in incomplete code due to the lack of inputs, making it difficult for Claude to reason about code execution and runtime behaviors.
The significant enhancement from CodeFlow is particularly valuable in incomplete code, where direct execution is infeasible due to the nature of incomplete code. By using Claude to generate likely inputs and CodeFlow to predict runtime errors without requiring external library setups, we effectively address this challenge. This integrated approach not only increases the detection rate of runtime errors but also provides reasonably accurate localization of the specific lines causing these errors. Other advanced LLM-based fuzzers, such as Fuzz4All [20], could be used with our framework.
@@ XIII Related Work
Early neural network-based fault localization (FL) approaches [23, 24, 25, 26, 27] primarily rely on test coverage data and struggle to distinguish between elements executed by failed tests and actual faulty elements [27]. In contrast, recent deep learning-based approaches such as GRACE [28], DeepFL [29], CNNFL [30], and DeepRL4FL [31] have demonstrated improved performance. GRACE introduces a novel graph representation for methods and learns to rank faulty methods.
@@ XIV Conclusion & Future Work
Current code models often overlook dynamic dependencies between lines of code, focusing only on plain text or correct execution traces. To address this, we introduce CodeFlow, an approach that predicts code coverage and detects runtime errors by learning both static and dynamic dependencies. CodeFlow uses CFGs and a Gated recurrent unit (GRU) network to represent execution paths and learn vector embeddings for CFG nodes. It also leverages execution traces via CFG to capture dynamic dependencies among statements. Our evaluation shows that CodeFlow significantly improves coverage prediction accuracy and runtime error localization, outperforming state-of-the-art models.
While CodeFlow has shown promising results, there are several avenues for future research and improvements:
  • Extending to other programming languages: Currently, CodeFlow has been tested primarily on Python code. Future work could involve adapting the model to support other popular programming languages, expanding its applicability.
  • Integration with development environments: Exploring ways to integrate CodeFlow into integrated development environments (IDEs) could provide real-time code coverage and error detection capabilities to developers.
  • Handling larger codebases: Investigating techniques to handle larger, more complex codebases and entire software projects would enhance its practical utility.
  • Combining with LLMs: Exploring hybrid approaches that leverage the strengths of both CodeFlow and LLMs could potentially lead to even more powerful code analysis tools.
  • Dynamic code analysis: Extending CodeFlow to perform more complex dynamic analysis tasks, such as predicting performance bottlenecks or potential vulnerabilities.
  • Continuous learning: Developing mechanisms for CodeFlow to continuously learn and improve its predictions based on feedback could enhance its effectiveness.
In conclusion, CodeFlow represents a significant step forward in predicting runtime program behavior without execution. By effectively modeling both static and dynamic code dependencies, it opens up new possibilities for code analysis and software quality assurance. As we continue to refine and expand this approach, we anticipate CodeFlow and similar graph-learning-based models will play an increasingly important role in the future of the static prediction of runtime behaviors. Our data and code is available in our website [39].

### v3

@@ XI-A Data Processing and Evaluation Metrics
In this section, we focus on a new metric, Error Localization Accuracy (ELA), which measures the number of times the predicted buggy line matches the actual buggy line. Based on the results from Section X, we observed that Claude performed the best in detecting runtime errors. Therefore, for this experiment, we use Claude as the main baseline to compare with our model.
@@ XI-B1 HeatMap Visualization
Alpha Value ELA (%)
$\alpha=0.5$ 72.22
$\alpha=0.7$ 74.82
$\alpha=0.9$ 77.40
$\alpha=0.95$ 78.31
@@ XII-A Fuzz Testing Procedure
  1. 1.
    Input Generation: We used Claude to generate inputs that are likely to raise runtime errors in the provided snippets.
  2. 2.
    Runtime Error Detection: The code snippet with the generated inputs was fed into CodeFlow. The model processed the code to determine whether it contained a runtime error and, if so, localized the buggy statement.
  3. 3.
@@ XII-B Empirical Results
ELA (%) 42.27 49.53 50.00 42.61
To evaluate CodeFlow’s effectiveness in supporting fuzz testing, we tested 50 buggy Python snippets from the FixEval dataset after removing all input variables and import statements. The results (Table VI) show that Claude alone detected 32 runtime errors, while its integration with CodeFlow detected 44 errors in 30 seconds (a 37.5% improvement), 46 errors in 60 seconds, and 47 errors in 120 seconds.
Additionally, the Error Localization Accuracy (ELA) with CodeFlow was consistently higher than that of Claude, as showed in RQ3. It is particularly challenging for LLMs like Claude to detect the correct line containing bugs in incomplete code due to the lack of inputs, making it difficult for Claude to reason about code execution and runtime behaviors.
The significant enhancement from our model is especially valuable for incomplete code, where direct execution is infeasible. By combining Claude to generate inputs and CodeFlow to predict runtime errors without external library setups, this approach effectively addresses the challenge. It not only improves runtime error detection rates but also provides accurate fault localization. LLM-based fuzzers, e.g., Fuzz4All [25] could also seamlessly integrate in our framework.
@@ XIII Related Work
Early ML fault localization approaches [28, 29, 30, 31, 32] primarily rely on test coverage data and struggle to distinguish between elements executed by failed tests and actual faulty elements [32]. In contrast, recent deep learning-based approaches such as GRACE [33], DeepFL [34], CNNFL [35], and DeepRL4FL [36] have demonstrated improved performance. GRACE introduces a novel graph representation for methods and learns to rank faulty methods.
@@ XIV Conclusion & Future Work
Current code models often overlook dynamic dependencies between lines of code, focusing only on plain text or correct execution traces. To address this, we introduce CodeFlow, an approach that predicts code coverage and detects runtime errors by learning both static and dynamic dependencies. CodeFlow uses CFGs and a GRU network to represent execution paths and learn vector embeddings for CFG nodes. It also leverages execution traces via CFG to capture dynamic dependencies among statements. Our evaluation shows that CodeFlow significantly improves coverage prediction accuracy and runtime error localization, outperforming state-of-the-art models. Our data and code is available in our website [44].
Acknowledgment
Tien N. Nguyen was supported in part by the US National Science Foundation (NSF) grant CNS-2120386 and the National Security Agency (NSA) grant NCAE-C-002-2021.